OpenStudy (anonymous):

If u and v are the vectors below, find the vector w whose tail is at the point halfway from the tip of v to the tip of u+v and whose head is at the point halfway from the tip of u to the tip of u−v. Assume all vectors are in standard position. u:(-3,-4,-4), v:(-5,-3,3). Answer is w=(6,2.5,-6.5)

OpenStudy (chosenmatt):

@WolfMaster

OpenStudy (anonymous):

The steps i've taken are: Find u+v=(-8,-7,-1) Find u-v=(2,-1,-7)